Subtract and write in simplest form:
12 10
---- - ----
Z Z
Is the answer:
2
---
Z^2
Thank you very much

Answer by checkley71(8403)

THE MISTAKE YOU MADE IS TO SUBTRACT THE NUMERATORS AND THEN MULTIPLIED THE DENOMINATORS.
12/Z-10/Z=(12-10)/Z=2/Z THE DENOMINATORS ARE EQUAL THEREFORE YOU JUST ADD/SUBTRACT THE NUMERATORS.
EXAMPLES:
1/4+2/4=(1+2)/4=3/4
5/6-2/6=(5-2)/6=3/6=1/2
8/15-4/15=(8-4)/15=4/15
---------------------------------------------
IF YOU ARE MULTIPLYING FRACTIONS THEN JUST MULTIPLY THE NUMERATORS & THE DENOMINATORS THUS:
12/Z*10/Z=120/Z*Z=120/Z^2
1/4*2/4=2/16=1/8
5/6*-2/6=-10/36=-5/18
8/15*-4/15=-32/225